OpenStudy (sarahezzmck):

What is the oblique asymptote of the function f(x) = the quantity of x squared plus 6 x plus 10, all over x plus 2? Will someone please help me with this problem?

OpenStudy (sarahezzmck):

@kewlgeek555 ??

OpenStudy (sarahezzmck):

y = x + 1 y = x - 1 y = x + 4 y = x - 4

OpenStudy (sarahezzmck):

@campbell_st ?

OpenStudy (sarahezzmck):

@Isaiah.Feynman

OpenStudy (phi):

It is difficult to read equations typed out as words. Use the equation editor (see button at the bottom of the input area) \[ f(x) = \frac{x^2+6x+10}{x+2}\]

OpenStudy (sarahezzmck):

oh I forgot about that but yes that is the equation. I'm sorry, I just totally forgot :(

OpenStudy (phi):

to find the asymptotes, divide x+2 into x^2 +6x + 10 Can you do that ? what do you get ?

OpenStudy (sarahezzmck):

\[x ^{3}+3x+5\] ?

OpenStudy (sarahezzmck):

Wait theres no x in 3x its just 3

OpenStudy (phi):

when you divide, you get a smaller polynomial See this https://www.khanacademy.org/math/algebra/multiplying-factoring-expression/dividing_polynomials/v/polynomial-division for how to do it.

OpenStudy (sarahezzmck):

ok thank you :)

OpenStudy (phi):

you should get